OpenStudy (anonymous):

A manufacturer of bird feeders cuts spacers from a tube that is 6 3/5 inches long. How many spacers can be cut from the tube if each spacer must be 5/7 inch thick? estimated? exact?

OpenStudy (anonymous):

6 3/5 divided by 5/7 convert 6 3/5 to improper fraction, invert 5/7, and multiply

OpenStudy (anonymous):

i got 9/

OpenStudy (anonymous):

I mean 9, I got 9. Am i correct ?

OpenStudy (anonymous):

33/ 5 * 7/5 = 231/25 = 9

OpenStudy (anonymous):

yup 9 with a remainder 6

OpenStudy (anonymous):

which means there is a small tube of thickness 6/25 left over

OpenStudy (anonymous):

thank you so much .. i am getting better with fractions !!